OpenStudy (mar1977):

Solve the equation. A. –39.71 B. –26.29 C. 39.71 D. 26. can really need a answer

OpenStudy (mar1977):

33=p-6.71 is the problem

OpenStudy (anonymous):

Let's solve your equation step-by-step. 33=p−6.71 Step 1: Flip the equation. p−6.71=33 Step 2: Add 6.71 to both sides. p−6.71+6.71=33+6.71 p=39.71

OpenStudy (anonymous):

Do you understand?

OpenStudy (mar1977):

yes thank it help to know I did the problem right thank very much